OpenStudy (anonymous):

Find the equation of the line through points (2, 5) and (4, 0).

OpenStudy (anonymous):

slope=-5/2 ,5x+2y-20=0

OpenStudy (lalaly):

first find slope \[\large{=\frac{y_2-y_1}{x_2-x_1}}=\frac{0-5}{4-2}=\frac{-5}{2}\] \[y-5=\frac{-5}{2}(x-2)\]\[y-5=\frac{-5}{2}+5\]\[y=\frac{-5}{2}x+10\]